OpenStudy (anonymous):

What are the solutions of the inequality? 12x – 3x + 11 > 4x – (17 – 9x) A. x > –7 B. x < 7 C.x < - 14/11 D.x > -14/11

OpenStudy (merchandize):

x The problem to solve is: 12x - 3x + 11›4x - ( 17-9x) First, let's work on the left hand side of your inequality, the 12x - 3x + 11 This means, for instance, to see if it can be simplified at all. Multiply x and 12 Multiply x and 1 The x just gets copied along. The answer is x x 12*x evaluates to 12x Multiply x and 3 Multiply x and 1 The x just gets copied along. The answer is x x 3*x evaluates to 3x 12*x-3*x evaluates to 9x 12*x-3*x+11 evaluates to 9x+11 So, all-in-all, the left hand side of your inequality can be written as: 9x+11 Now, let's work on the right hand side of your inequality, the 4x - ( 17-9x) Multiply x and 4 Multiply x and 1 The x just gets copied along. The answer is x x 4*x evaluates to 4x Multiply x and 9 Multiply x and 1 The x just gets copied along. The answer is x x 9*x evaluates to 9x 17-9*x evaluates to 17-9x 4x - -9x = 13x The answer is 13x-17 4*x-(17-9*x) evaluates to 13x-17 The right hand side of your inequality can be written as: 13x-17 So with these (any) simplifications, the inequality we'll set out to solve is: 9x+11 › 13x-17 Move the 11 to the right hand side by subtracting 11 from both sides, like this: From the left hand side: 11 - 11 = 0 The answer is 9x From the right hand side: -17 - 11 = -28 The answer is -28+13x Now, the inequality reads: 9x › -28+13x Move the 13x to the left hand side by subtracting 13x from both sides, like this: From the left hand side: 9x - 13x = -4x The answer is -4x From the right hand side: 13x - 13x = 0 The answer is -28 Now, the inequality reads: -4x › -28 To isolate the x, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the x on the left side of the inequality. The last step is to divide both sides of the inequality by -4 like this: To divide x by 1 The x just gets copied along in the numerator. The answer is x -4x ÷ -4 = x -28 ÷ -4 = 7 Note! Here's an important point about inequalities: We are now dividing by -4, which is a number less than zero! This changes the form of your inequality. Or, in other words, your › will now become a ‹. The solution to your inequality is: x ‹ 7 So, your solution is: x must be less than 7